# A Scoping Review of Synthetic Data Generation for Biomedical Research and Applications

### v1

@@ Abstract
Synthetic data generation––mitigating data scarcity, privacy concerns, and data quality challenges in biomedical fields––has been facilitated by rapid advances of large language models (LLMs). This scoping review follows PRISMA-ScR guidelines and synthesizes 59 studies, published between 2020 and 2025 and collected from PubMed, ACM, Web of Science, and Google Scholar. The review systematically examines biomedical research and application trends in synthetic data generation, emphasizing clinical applications, methodologies, and evaluations. Our analysis identifies data modalities of unstructured texts (78.0%), tabular data (13.6%), and multimodal sources (8.4%); generation methods of prompting (72.9%), fine-tuning (22.0%) LLMs and specialized model (5.1%); and heterogeneous evaluations of intrinsic metrics (27.1%), human-in-the-loop assessments (55.9%), and LLM-based evaluations (13.6%). The analysis addresses current limitations in what, where, and how health professionals can leverage synthetic data generation for biomedical domains. Our review also highlights challenges in adaption across clinical domains, resource and model accessibility, and evaluation standardizations.
Introduction
@@
Data is the key to train reliable AI models for broad biomedical and clinical applications, such as medical diagnosis [1, 2], therapeutic treatments [3], and drug discovery [4]. However, obtaining massive, privacy-preserving, and high-quality data faces critical challenges, such as data availability, noisy and missing data, and legal regulations. Increasing biomedical studies are turning attentions to synthetic data generation, a process of creating artificial datasets that accurately replicate the statistical and structural properties of real-world data. Nonetheless, creating high-quality synthetic data remains challenging due to the inherent heterogeneity, complexity, and variability characteristic of biomedical data. Large language models (LLMs), as a generative AI method, offer a promising solution by enhancing the accuracy and diversity of synthetic datasets through advanced learning algorithms.
@@
Synthetic data generation in the biomedical and clinical fields has achieved significant advances by LLMs (e.g., GPT-4 [5] and Llama 3 [6]) in the recent years, as shown in Figure 2. For example, LLMs have been applied successfully to obtain clinical narratives and simulate real patient records of mental health for diagnosis and behavior analysis [7], such as predicting suicide [8] and depression patterns [9]. The strong generation capability not only relieve data privacy concerns [10] to facilitate model developments but also may promote data quality [11] to support clinical decision-making. For example, synthetic text corpora have augmented classification models for cardiovascular and Alzheimer’s disease diagnosis [12]; A study develops a multi-agent dialogue generation framework (NoteChat) that creates synthetic patient-physician conversations to improve clinical documentation [13]. Nevertheless, questions and uncertainties remain regarding the best practices, validation approaches, and specific application areas for those LLM-generated synthetic data.
@@
The goal of this scoping review is to comprehensively summarize and assess recent research publications on the biomedical and clinical applications leveraging LLMs for synthetic data generation. While multiple reviewing studies [14, 15, 16] have covered the LLM-generated synthetic data for general domains, studies on biomedical and clinical domains are still not available. For example, a close study [16] examines recent developments of generation algorithms and models on news and social media domains instead of biomedical fields. Specifically, through this study, we will identify and present the current state-of-the-art models, technical methods, evaluations or assessments of synthetic data quality, and gaps and opportunities for future research. We seek to answer a concrete yet unsolved question: what biomedical and clinical applications can be effectively addressed using LLMs-generated synthetic data, and by how?
Results
@@
Our search and initial review resulted in 132 articles and finalized with 59 articles for this scoping review. XH, HW, and HR reviewed full texts of the articles and removed 73 articles due to ineligible article types (n=2), no large language model (n=7), unrelated to biomedical data generation (n=18), and not peer-reviewed and published in a journal or conference proceedings (n=46). Fifty-nine articles met our criteria and were included in our subsequent analyses. In terms of publication venues and formats, the included literature consists of 43 journal articles (72.9%) and 16 peer-reviewed conference papers (27.1%), leaving 59 studies to be included in this scoping review. Figure 1 shows that the study number of biomedical synthetic data generation climb steadily over years. Particularly, the substantial growth after 2022 reflects a particularly evolving trend of synthetic biomedical data generation via LLM, an emerging generation tool in recent years.

### v2

@@ Abstract
Synthetic data generation using large language models (LLMs) demonstrates substantial promise in addressing biomedical data challenges and shows increasing adoption in biomedical research. This study systematically reviews recent advances in synthetic data generation for biomedical applications and clinical research, focusing on how LLMs address data scarcity, utility, and quality issues with different modalities. We conducted a scoping review following PRISMA-ScR guidelines and searched literature published between 2020 and 2025 through PubMed, ACM, Web of Science, and Google Scholar. A total of 59 studies were included based on relevance to synthetic data generation in biomedical contexts. Among the reviewed studies, the predominant data modalities were unstructured texts (78.0%), tabular data (13.6%), and multimodal sources (8.4%). Common generation methods included LLM prompting (74.6%), fine-tuning (20.3%), and specialized models (5.1%). Evaluations were heterogeneous: intrinsic metrics (27.1%), human-in-the-loop assessments (44.1%), and LLM-based evaluations (13.6%). However, limitations and key barriers persist in data modalities, domain utility, resource and model accessibility, and standardized evaluation protocols. Future efforts may focus on developing standardized, transparent evaluation frameworks and expanding accessibility to support effective applications in biomedical research.
keywords:
synthetic data generation, language model, biomedical informatics, data quality
1 Introduction
@@
Data is the key to train reliable AI models for broad biomedical and clinical applications, such as medical diagnosis [liu2025time, jones2025examining], therapeutic treatments [stade2024large], and drug discovery [blanco2023role]. However, obtaining massive, privacy-preserving, and high-quality data faces critical challenges, such as data availability, noisy and missing data, and legal regulations. Increasing biomedical studies are turning attention to synthetic data generation, a process of creating artificial datasets that accurately replicate the statistical and structural properties of real-world data. Nonetheless, creating high-quality synthetic data remains challenging due to the inherent heterogeneity, complexity, and variability characteristic of biomedical data. Large language models (LLMs) may offer a promising solution by enhancing the availability and utility of synthetic datasets in critical applications, such as clinical note analysis for stroke thrombolysis contraindication identification [chen2025automated], mental health diagnosis through automated interview assessment [wu2024callm], and radiology report generation for fracture misdiagnosis detection [liu2025generating].
@@
Synthetic data generation in biomedical and clinical fields has achieved significant advances by LLMs (e.g., GPT-4 [achiam2023gpt] and Llama 3 [dubey2024llama]) in recent years, as shown in Figure 1 and 5. For example, LLMs have been applied successfully to obtain clinical narratives and simulate real patient records of mental health for diagnosis and behavior analysis [han2024chain], such as predicting suicide [ghanadian2024socially] and depression patterns [bucur2024leveraging]. The strong generation capability not only relieves data privacy concerns [sarkar2024identification] to facilitate model developments but also may promote data quality [xu-etal-2024-knowledge] to support clinical decision-making. For example, synthetic corpora have augmented classification models for cardiovascular and Alzheimer’s disease diagnosis [li2023two]; and a study develops a multi-agent dialogue generator (NoteChat) that creates synthetic patient-physician conversations to improve clinical documentation [wang2024notechat]. Nevertheless, questions and uncertainties remain regarding the best practices, validation approaches, and specific application areas for those LLM-generated synthetic data.
@@
The goal of this scoping review is to comprehensively summarize and assess recent research publications on the biomedical and clinical utilities leveraging LLMs for synthetic data generation and its quality evaluation. While multiple reviewing studies [figueira2022survey, long-etal-2024-llms, li-etal-2023-synthetic] have covered LLM-generated synthetic data for general domains, studies on biomedical and clinical domains remain underexplored. For example, a close study [li-etal-2023-synthetic] examines recent developments of generation algorithms and models on news and social media domains instead of biomedical fields. Several recent review studies [murtaza2023synthetic, gonzales2023synthetic, smolyak2024large, loni2025review] have examined the topic in health and clinical fields, however, their focus on methodologies and models often overlooks fundamental aspects of synthetic data itself, such as types, utility, accessibility, and data quality —– gaps that this study aims to address. Specifically, through this study, we will identify and present the current state-of-the-art models, technical methods, evaluations or assessments of synthetic data quality, and gaps and opportunities for future research. We seek to answer a concrete yet unsolved question: what biomedical and clinical applications can be effectively addressed using LLM-generated synthetic data, and how?
2 Results
@@
Our search and initial review resulted in 132 articles and finalized with 59 articles for this scoping review. XH, HW, and HR reviewed full texts of the articles and removed 73 articles due to ineligible article types ($n=2$), no large language model ($n=7$), unrelated to biomedical data generation ($n=18$), and not peer-reviewed and published in a journal or conference proceedings ($n=46$). Inter-rater reliability, assessed using Cohen’s $\kappa$ [landis1977measurement], indicated almost perfect agreement for both screening phases ($\kappa=0.96$ for title/abstract screening, $n=307$; $\kappa=0.95$ for full-text screening, $n=132$). Fifty-nine articles met our criteria and were included in our subsequent analyses. In terms of publication venues and formats, the included literature consists of 43 journal articles (72.9%) and 16 peer-reviewed conference papers (27.1%), leaving 59 studies to be included in this scoping review. Figure 1 shows that the study number of biomedical synthetic data generation climbs steadily over the years. Particularly, the substantial growth after 2022 reflects a particularly evolving trend of synthetic biomedical data generation via LLM, an emerging generation tool in recent years.